An insurance company wants to offer a policy where it pays out $50,000 for death and $20,000 for disability. It calculates a probability of .0001 that a policy holder will die and a probability of .0003 that a policy holder will become disabled during the course of a year. The company wants to have an expected profit of $25 per year per policyholder. What should the company charge for its annual premium?
Answer #1

Profit = Sales- Expense................1

Let the Company charge X as Annual Premium of 1 policy

The Probable Expense = 50,000*0.0001+ 20000*0.0003= 5+6=11

From Eq.1 ... Sales= Profit+Expense= 25+11

Therefore the Company Should Charge 36$ as Annual Premiuim